OpenStudy (anonymous):

A car is traveling at a rate of 25 meters per second. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 2 hours? Do not round your answers

OpenStudy (neer2890):

first convert 25 m into kilometer. 25m=25/1000km=1/40km and then convert second into minutes i.e., 1sec=1/3600 minutes

OpenStudy (dumbcow):

\[25 \frac{m}{s} * \frac{1 km}{1000 m}*\frac{3600 s}{1 hr}\]

OpenStudy (neer2890):

25m/sec=(1/40)/(1/3600)=3600/40=90km/hr